Equipment and Calibration History
Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 [62] 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
HMP45AC T1250012 | 2013-04-19 | Lab none | Installed | Tonzi-Tower | Installed on the Tonzi tower at the 2nd from bottom (1.8m) CO2 profile inlet location. Wired to the AM16/32 multiplexor connected to the tower top CR23X |
HMP45AC T0720012 | 2013-04-19 | Lab none | Installed | Tonzi-Tower | Installed on the Tonzi tower at the bottom (0.35m) CO2 profile inlet location. Wired to the AM16/32 multiplexor connected to the tower top CR23X |
HFP01SC 00067 | 2013-04-24 | Lab none | Installed | Tonzi-SoilCO2 | Installed as SHF#2 on the soil CO2 CR23X and placed in the ground to the south of the data logger box |
HFP01 000141 | 2013-04-24 | Lab none | Installed | Tonzi-SoilCO2 | Installed as SHF#1 on the soil CO2 CR23X and placed in the ground to the south of the data logger box |
LI-6262 IRG3-1079 | 2013-04-22 | Lab none | Calibration | Lab none | Calibrated in differential mode for check of new secondary standard. Ambient temperature was 28.8C, pressure was 100.55kPa. Using N2 (1100psi) and god bottle 402.69ppm CO2 (1800psi).
Flowing N2 in both sample and reference CO2 +3ppm - adjusted to 0.0 H2O -0.19 - railed Flowing N2 in reference and god bottle in sample CO2 = 401.5ppm adjusted to 402.69ppm. Flowing N2 in reference and new secondary standard in sample, I recorded serial data with Hyperterm and took an average of 587 points. The new tank is 404.02ppm |
LI-7500A 0035 | 2013-04-19 | Tonzi-Tower | Removed | Lab none | Removed from the Tower for calibration. Replaced by sn 0418 |
LI-7500A 0042 | 2013-04-19 | Tonzi-Floor | Removed | Lab none | Removed from the Floor tower for calibration. Replaced by sn 1594 |
LI-7500A 0418 | 2013-04-19 | Lab none | Installed | Tonzi-Tower | Swapped on to the Tower to replace sn 0035 for calibration |
LI-7500A 1594 | 2013-04-19 | Lab none | Installed | Tonzi-Floor | Swapped on to the Floor Tower to replace sn 0042 for calibration. |
LI-7500A 1594 | 2013-04-18 | Lab none | Calibration | Lab none | Last calibration: 2012-10-23 AGC=51 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.52kPa, Tair=23.81C, Tdew=6.77C Zero Air at 1700psi, CO2 in Air 392.3 at 1700 Tdewspan=15.04 CO2 zero 0.8674 -> 0.8649 Offset: -9.87ppm CO2 span 1.0021 -> 1.0179 Offset: -7.8ppm H2O zero 0.8857 -> 0.8963 Offset: 56.4ppm H2O span 0.8959 -> 0.9956 Offset: -2.18ppm Notes: Offsets a little larger than usual |
LI-7500A 0418 | 2013-04-18 | Lab none | Calibration | Lab none | Last calibration: 2012-10-23 AGC=43 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.71kPa, Tair=22.02C, Tdew=4.42C Zero air at 1700psi, CO2 in air 392.30 at 1700 Tdewspan=15.04 CO2 zero 0.9011 -> 0.9017 Offset: 2.26ppm CO2 span 1.0085 -> 1.0104 Offset: -0.9ppm H2O zero 0.8516 -> 0.8497 Offset: -10.7ppm H2O span 1 -> 0.9871 Offset: ppm Notes: |
GMP222 H1630066 | 2013-04-04 | Lab none | Maintenance | Lab none | |
GMP222 H1630066 | 2013-03-28 | Mayberry CO2 Probe B | Removed | Lab none | We removed the probe since it was reading 5000mV even we were put the sensor in the air. |
DO6400/T DO-07 | 2013-03-14 | Lab none | Installed | Mayberry DO Sensor B | Reinstalled lower DO probe at the Tule tower |
LI-7700 TG1-0155 | 2013-03-20 | Lab none | Calibration | Lab none | Calibration Location: Lab Last calibration: 2012-10-31 RSSI=89%, Optical RH=4% Ambient P=100.8kPa, Tair=23C UHP N2 at 1100psi, 2nd CH4 in air 3.875 at 1800 CH4 Zero Offset: 0.0ppm CH4 Span Offset: 0.035ppm Notes: Last calibrated by Frank Anderson. |
LI-7500A 0418 | 2013-02-27 | Vaira Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Vaira Tower by Siyan for calibration. Replaced by sn 0976. |
LI-7500 0976 | 2013-02-27 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed on the Vaira tower by Siyan to replace sn 0418 for calibration. |
LI-7500A 75H-2182 | 2013-02-14 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed to replace sn 2150 for upgrade at Licor |
LI-7550 AIU-0397 | 2013-02-14 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed to replace sn 0614 for upgrade at Licor |
LI-7500A 75H-2150 | 2013-02-14 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Mayberry tower with AIU-0614 for upgrade at Licor |
LI-7550 AIU-0614 | 2013-02-14 | Mayberry Tower | Removed | Lab none | Removed from the Mayberry tower with head 2150 for upgrade at Licor |
FMA RMT-200 08-0114 | 2013-03-14 | Lab none | Installed | Sherman Island | Installed in the Sherman Island box by Cove to replace sn 05-0010. 05-0010 was removed due to low pressure in low flow mode. |
FMA RMT-200 05-0010 | 2013-03-14 | Sherman Island | Removed | Lab none | Removed from the Sherman Island box by Cove due to low pressure in the low flow mode. Replaced by sn 08-0114. |
FMA RMT-200 05-0010 | 2013-03-18 | Lab none | Calibration | Lab none | Last calibration: 2012-08-23 Cell Pressure=157.32Torr, Cell Temperature=24.40C, Ring Time=12.694usec 2nd CH4 in Air at psi CH4 Span=3.875ppm CH4 = 3.888 -> 3.872 Offset: -13ppb Notes: Done in the lab after cleaning the 10um filter infront of the pressure controller |
AM16/32A 1237 | 2013-03-13 | Lab none | Installed | Tonzi-Tower | Installed in the cr23x box near the top of the tower in preparation for adding new diffuse PAR and 4-band radiometer. Currnetly not connected. |
DO6400/T DO-07 | 2013-03-12 | Lab none | Calibration | Lab none | DO-07 was calibrated in the lab: M = 8.73/36.8 (mg/L)/mV (aka ppm/mV) @ temp = 21.5 C and press.= 101.5 Kpa |
DO6400/T DO-07 | 2013-02-28 | Lab none | Removed | Lab none | Removed for testing in lab since when holding it out of water at Mayberry the readings were decreasing instead of going back to atmospheric. |
DO6400/T DO-07 | 2013-03-12 | Lab none | Maintenance | Lab none | Replaced membrane. Previous one appeared to be torn. readings seem much better now. |
DO6400/T DO-04 | 2013-02-28 | Lab none | Installed | Mayberry DO Sensor A | Installed in the surface at Mayberry about 10 cm below surface. |
DO6400/T DO-07 | 2013-02-28 | Mayberry DO Sensor B | Removed | Lab none | Removed for testing in lab since when holding it out of water at Mayberry the readings were decreasing instead of going back to atmospheric. |
FMA RMT-200 08-0114 | 2013-03-12 | Lab none | Maintenance | Lab none | The ring time was around 11.8usec so I cleaned the mirrors before likely putting it in at Sherman later this week. After two attempts at cleaning the mirrors the ring time seems to be steady at about 16.8usec |
DO6400/T DO-04 | 2013-02-26 | Lab none | Maintenance | Lab none | This sensor was pulled from the rice since the field was drained. I (Sara) cleaned it and replaced the membrane since it had been a while. I left it without electrolyte solution until we install it in the field again to minimize corrosion. |
LI-7500 0976 | 2013-02-26 | Lab none | Calibration | Lab none | Last calibration: 2012-11-06 AGC=58 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.10kPa, Tair=16.57C, Tdew=7.62C Zero air at 1600psi, 2nd CO2 in air 392.30 at 1800 Tdewspan=15.05 CO2 zero 0.8599 -> 0.8594 Offset: -1.61ppm CO2 span 1.0056 -> 1.0105 Offset: -2.45ppm H2O zero 0.6233 -> 0.6290 Offset: 42.2ppm H2O span 1.0086 -> 1.0101 Offset: -0.03ppm Notes: H2O span fudged by 0.07C |
DO6400/T DO-06 | 2013-02-25 | Lab none | Maintenance | Lab none | This sensor was givning low readings. Sara cleaned, replaced the membrane and refilled the electrolyte for this sensor. Afterwards its readings were reasonable. |
DO6400/T DO-06 | 2013-02-21 | Mayberry DO Sensor A | Removed | Lab none | Removed from the upper water location near the tule tower at Mayberry. It was reading low and did not climb when exposed to the air. Brought to the lab for maintenance. |
DO6400/T DO-04 | 2013-02-21 | Twitchell Rice DO Sensor | Removed | Lab none | Removed from the now drained field at Tiwtchell. There will be no water in the field until after the rice is up late this coming spring. |
LI-7700 TG1-0211 | 2013-02-21 | Twitchell Rice Gas Analyzer CH4 | Calibration | Twitchell Rice Gas Analyzer CH4 | Calibration Location: Twitchell Last calibration: 2012-09-06 RSSI=65%, Optical RH=6% Ambient P=102.1kPa, Tair=7.87C Zero Air at 1800psi, 2nd CH4 in Air 4.982 at 1300 CH4 Zero Offset: 0.005ppm CH4 Span Offset: 0.018ppm Notes: |
LI-7700 TG1-0224 | 2013-02-21 | Mayberry Gas Analyzer CH4 | Calibration | Mayberry Gas Analyzer CH4 | Calibration Location: Mayberry Last calibration: 2012-09-06 RSSI=85%, Optical RH=3% Ambient P=101.97kPa, Tair=14.4C Zero Air at 1800psi, 2nd CH4 in Air 4.982 at 1300 CH4 Zero Offset: 0.011ppm CH4 Span Offset: 0.010ppm Notes: The 7700 had lost line lock and was manually lock prior to calibration. |
LI-7700 TG1-0223 | 2013-02-19 | West Pond Gas Analyzer CH4 | Calibration | West Pond Gas Analyzer CH4 | Calibration Location: West Pond Last calibration: 2012-09-06 RSSI=66.8%, Optical RH=1.0% Ambient P=101.98kPa, Tair=10.8C Zero Air at 1800psi, 2nd CH4 in Air 4.982 at 1300 CH4 Zero Offset: 0.016ppm CH4 Span Offset: 0.006ppm Notes: |
LI-7550 AIU-0674 | 2013-02-12 | Factory | Upgrade | Lab none | Returned from Licor with 75H-2176 after repair for the condensation problem. The external connector circuit board at the bottom of the box has been completely sealed. New firmware v6.0.0 has been loaded on the box and requires new Windows interface software. Pressure Sensor sn PX-2828 A0=57.462, A1=15.200. |
LI-7500A 75H-2176 | 2013-02-12 | Lab none | Upgrade | Lab none | Returned from Licor with AIU-0674 after repair of condensation problem. The firmware was updated and it was recalibrated.
CO2: A=1.50515E2, B=4.94461E2, C=4.84055E7, D=-1.58118E10, E=2.05944E12, XS=-0.0014, Z=-2.10000E-3, zero=0.8676, span=1.0038 H2O: A=4.89101E3, B=3.36421E6, C=-5.74265E7, XS=-0.0014, Z=2.13000E-2, zero=0.8539, span=0.9906, CX=40300 SS: B=2.213, C=3.942 |
LI-7550 AIU-0397 | 2013-02-11 | Factory | Upgrade | Lab none | Returned from Licor with 75H-2182 after repair for the condensation problem. The external connector circuit board at the bottom of the box has been completely sealed. New firmware v6.0.0 has been loaded on the box and requires new Windows interface software. Pressure Sensor sn PX-2379 A0=57.091, A1=15.210. |
LI-7500A 75H-2182 | 2013-02-11 | Factory | Upgrade | Lab none | Returned from Licor with AIU-0397 after repair of condensation problem. The firmware was updated and it was recalibrated.
CO2: A=1.59506E2, B=1.38372E3, C=5.57490E7, D=--1.93676E10, E=2.63113E12, XS=-0.0002, Z=-4.20000E-3, zero=0.8837, span=1.0037 H2O: A=5.37667E3, B=4.39566E6, C=-3.30591E8, XS=-0.0013, Z=1.62000E-2, zero=0.8417, span=0.9978, CX=41300 SS: B=2.154, C=4.196 |
FMA RMT-200 08-0114 | 2013-02-12 | Lab none | Calibration | Lab none | Last calibration: 2013-01-29 Cell Pressure=143.19Torr, Cell Temperature=22.14C, Ring Time=11.907usec 2nd CH4 in air at 1400psi CH4 Span=4.982ppm CH4 = 4.960 -> 4.982 Offset: 22ppb Notes: Calibration for check of new 400ppm CO2 cal tank. After calibrating the LGR data was recorded from the new tank. An average of 146 point was 3.875ppm CH4 |
LI-6262 IRG3-1079 | 2013-02-12 | Lab none | Calibration | Lab none | Calibrated in differential mode for check of new secondary standard. Ambient temperature was 23.55C, pressure was 101.98kPa. Using N2 (1600psi) and god bottle 402.69ppm CO2 (1800psi).
Flowing N2 in both sample and reference CO2 -3.79 - adjusted to 0.0 H2O -0.0863 - adjusted to 0.0 Flowing N2 in reference and god bottel in sample CO2 = 402.90ppm adjusted to 402.69ppm. Flowing N2 in reference and new secondary standard in sample, I recorded serial data with Hyperterm and took an average of 201 points. The new tank is 344.62ppm |
HP-127-1.0-1.3-71P P05 | 2013-01-16 | Lab none | Calibration | Lab none | Calibrated by Sebastian at University of British Columbia: Multiplier = 95.86(W/m2)/mV |
HP-127-1.0-1.3-71P P04 | 2013-01-16 | Lab none | Calibration | Lab none | Calibrated by Sebastian at University of British Columbia: Multiplier = 106.2(W/m2)/mV |
HP-127-1.0-1.3-71P P03 | 2013-01-16 | Lab none | Calibration | Lab none | Calibrated by Sebastian at University of British Columbia: Multiplier = 100.2(W/m2)/mV |
HP-127-1.0-1.3-71P P02 | 2013-01-16 | Lab none | Calibration | Lab none | Calibrated by Sebastian at University of British Columbia: Multiplier = 99.29(W/m2)/mV |
HP-127-1.0-1.3-71P P01 | 2013-01-16 | Lab none | Calibration | Lab none | Calibrated by Sebastian at University of British Columbia: Multiplier = 99.54 (W/m2)/mV |
HP-127-1.0-1.3-71P P05 | 2013-02-06 | Lab none | Installed | Vaira | Installed in a profile at -25cm to the east of the Vaira radiation tower |
HP-127-1.0-1.3-71P P04 | 2013-02-06 | Lab none | Installed | Vaira | Installed in a profile at -16cm to the east of the Vaira radiation tower |
HP-127-1.0-1.3-71P P03 | 2013-02-06 | Lab none | Installed | Vaira | Installed in a profile at -8cm to the east of the Vaira radiation tower |
HP-127-1.0-1.3-71P P02 | 2013-02-06 | Lab none | Installed | Vaira | Installed in a profile at -4cm to the east of the Vaira radiation tower |
HP-127-1.0-1.3-71P P01 | 2013-02-06 | Lab none | Installed | Vaira | Installed in a profile at -2cm to the east of the Vaira radiation tower |
GMP221 H0960006 | 2013-01-04 | Lab none | Tested | Disposed | Upon inspection in the lab. this sensor was found to be severely corroded. However, at the time of inspection the sensor was not damp. |
GMP221 H0960006 | 2013-01-03 | Mayberry CO2 Probe B | Removed | Lab none | Removed from the tules near the tule tower due to bad values. |
NL115 6761 | 2013-01-24 | Lab none | Installed | Vaira | Installed the the computer box with CR1000 sn 47108 to collect images from the Phenecam |
CR1000 47108 | 2013-01-24 | Lab none | Installed | Vaira | Installed the the computer box with NL115 sn 6761 to collect images from the Phenecam |
CFM100 4994 | 2012-08-22 | Lab none | Installed | Vaira | Installed with CR1000 sn 23865 to collect flux data at the Vaira tower in parallel with winflux. This data logger is in the box with the sonic PSU. |
CR1000 23865 | 2012-08-22 | Lab none | Installed | Vaira | Installed with NL115 sn 6761 to collect flux data at the Vaira tower in parallel with winflux. This data logger is in the box with the sonic PSU. |
GMP221 H1310008 | 2013-01-17 | Lab none | Installed | West Pond CO2 probe A | Installed in a teflon sock in the water near the other water sensors |
GMP222 H1630066 | 2013-01-17 | Lab none | Installed | Mayberry CO2 Probe B | Installed in a teflon sock in the water in the tule patch |
LI-7550 AIU-0674 | 2013-01-10 | Lab none | Shipped | Factory | Sent to Licor with H-2176 for fix of condensation problem |
LI-7500A 75H-2176 | 2013-01-10 | Lab none | Shipped | Factory | Sent to Licor with AIU-0674 for fix of condensation problem |
LI-7550 AIU-0397 | 2013-01-10 | Lab none | Shipped | Factory | Sent to Licor with H-2182 for fix of condensation problem |
LI-7500A 75H-2182 | 2013-01-10 | Lab none | Shipped | Factory | Sent to Licor with AIU-0379 for fix of condensation problem |
LI-7500A 0042 | 2013-02-06 | Lab none | Installed | Tonzi-Floor | Installed on the Floor Tower to replce 0976 for calibration |
LI-7500 0976 | 2013-02-06 | Tonzi-Floor | Removed | Lab none | Removed from the Floor tower for calibration. Replaced by 0042 |
LI-7500A 0035 | 2013-02-06 | Lab none | Installed | Tonzi-Tower | Installed on the Tonzi Tower to replace 1594 for calibration |
LI-7500A 1594 | 2013-02-06 | Tonzi-Tower | Removed | Lab none | Removed from the Tonzi Tower for calibration. Replaced by 0035. |
SATI-3K 000603 | 2013-01-22 | Lab none | Returned | Returned | Returned to Robin in Goldstein lab |
SATI-3K 000603 | 2012-12-11 | Tonzi-Tower | Removed | Lab none | Removed from the Tonzi tower - end of inter comparison |
SATI-3K 000603 | 2012-11-15 | Lab none | Installed | Tonzi-Tower | Installed on the Tonzi tower just to the east of the Gill sonic. This will be an inter-comparison of a few weeks. Data is being collected in separate tables on the CR1000 |
LI-7500A 0042 | 2013-02-04 | Lab none | Calibration | Lab none | Last calibration: 2012-08-02 AGC=52 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.2kPa, Tair=19.9C, Tdew=6.6C dry air at 1500psi, CO2 421.22 at 1700 Tdewspan=15.04 CO2 zero 0.9513 -> 0.9508 Offset: -1.84ppm CO2 span 1.0044 -> 1.0084 Offset: 1.57ppm H2O zero 0.8472 -> 0.8575 Offset: 0.54ppm H2O span 0.9696 -> 0.9872 Offset: 0.96ppm Notes: by Siyan |
LI-7500A 0035 | 2013-02-04 | Lab none | Calibration | Lab none | Last calibration: 2012-08-07 AGC=48 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.2kPa, Tair=18.9C, Tdew=6.6C dry air at 1500psi, CO2 421.22 at 1700 Tdewspan=15.04 CO2 zero 0.9388 -> 0.9417 Offset: 12.82ppm CO2 span 1.0176 -> 1.0150 Offset: 0.56ppm H2O zero 0.8994 -> 0.8846 Offset: -1.77ppm H2O span 0.9957 -> Offset: 4.24ppm Notes: all coefficients were zero before calibration. Joe helped to reset them successfully. |
LI-7500A 0041 | 2013-01-30 | Lab none | Installed | Sherman Island Gas Analyzer CO2/H2O | Installed on the Sherman tower to replace sn 0035 for calibration. |
LI-7500A 0035 | 2013-01-30 | Sherman Island Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Sherman tower for calibration. Replaced by sn 0041 |
LI-7500A 0073 | 2013-01-30 | Lab none | Installed | DWR Corn | Installed on the DWR-Corn tower to replace sn 0042 for calibration |
LI-7500A 0042 | 2013-01-30 | DWR Corn | Removed | Lab none | Removed from the DRW-Corn tower for calibration. Replaced with sn 0073 |
FMA RMT-200 08-0114 | 2013-01-29 | Lab none | Calibration | Lab none | Last calibration: 2012-8-29 Cell Pressure=142.76Torr, Cell Temperature=21.69C, Ring Time=11.939usec 2nd CH4 in air at 1400psi CH4 Span=4.982ppm CH4 = 4.959 -> 4.982 Offset: 23.000000000001ppb Notes: Calibrated to check new 400ppm CO2 secondary standard. I took and average of 75 points as recorded in the LGR internal files and got a concentration of 3.856ppm CH4 |
LI-7500A 0041 | 2013-01-29 | Lab none | Calibration | Lab none | Last calibration: 2012-07-24 AGC=52 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.65kPa, Tair=25.10C, Tdew=4.63C Zero air at 1700psi, CO2 in air 392.3 at 2000 Tdewspan=15.07 CO2 zero 1.0008 -> 1.0012 Offset: 1.35ppm CO2 span 1.0084 -> 1.0080 Offset: 0.3ppm H2O zero 0.8201 -> 0.8220 Offset: 10.8ppm H2O span 0.9947 -> 0.9980 Offset: 0.0ppm Notes: Very tight correspondence with last calibration |
LI-7500A 0073 | 2013-01-29 | Lab none | Calibration | Lab none | Last calibration: 2012-07-24 AGC=59 SDM=7 DAC1=0-1000 DAC2=10-25 Ambient P=101.58kPa, Tair=22.80C, Tdew=3.37C Zero air at 1700psi, CO2 in air 392.3 at 2000 Tdewspan=15.07 CO2 zero 0.9581 -> 0.9581 Offset: 0.10ppm CO2 span 1.0042 -> 1.0028 Offset: 0.6ppm H2O zero 0.7305 -> 0.7332 Offset: 16.4ppm H2O span 1.0154 -> 1.0186 Offset: 0.0ppm Notes: Very tight correspondence with last calibration |
LI-6262 IRG3-1079 | 2013-01-29 | Lab none | Calibration | Lab none | Calibrated in differential mode for check of new secondary standard. Ambient temperature was 22.9C, pressure was 101.84kPa. Using N2 (1600psi) and god bottle 402.69ppm CO2 (1800psi).
Flowing N2 in both sample and reference CO2 -0.53 - adjusted to 0.0 H2O +0.055 - railed but okay Flowing N2 in reference and god bottel in sample CO2 = 401.70ppm adjusted to 402.69ppm. Flowing N2 in reference and new secondary standard in sample, I recorded serial data with Hyperterm and took an average of 300 points. The new tank is 392.30ppm |
AM16/32A 1237 | 2012-09-06 | Tonzi | Removed | Lab none | Removed with all the above ground components of the C1 sap flow station. This sap flow station has not been working for a while and there are no plans to continue sap flow measurements in the future. |
CS450 20010026 | 2013-01-03 | Lab none | Installed | DWR Corn | Installed in the original well hole in the corn field. See fieldnotes of details. |
CS450 20010026 | 2012-12-18 | Lab none | Repair | Lab none | I pulled apart the existing splice in the cable, cut a small out and made a new splice. After wards I tested it in the lab and it seemed to function well. |
DO6400/T DO-04 | 2013-01-03 | West Pond 0 | Moved | Twitchell Rice DO Sensor | This sensor was moved from the upper location at West Pond to near the well at the Twitchell Rice. The sensor was left laying sideways on the bottom of the flooded field. |
LI-7550 AIU-0397 | 2013-01-03 | West Pond Interface LI-7550 | Removed | Lab none | Removed from the West Pond tower for repair of the condesation problem. Replaced by AIU-0666. Heads swapped for calibration at the same time. |
LI-7550 AIU-0397 | 2012-11-08 | Lab none | Installed | West Pond Interface LI-7550 | Installed on the West Pond tower to swap out AIU-650 for repair of the condensation problem. |
LI-7550 AIU-0666 | 2013-01-03 | Lab none | Installed | West Pond Interface LI-7550 | Installed on the West Pond tower with head 0065 to replace AIU-0397 for repair of condensation problem. |
LI-7550 AIU-0674 | 2013-01-03 | Twitchell Rice Interface LI-7500 | Removed | Lab none | Reomved from the Twitchell Rice tower for repair of the condensation problem, replaced by AIU-0650. Heads swapped for calibration at the same time. |
LI-7550 AIU-0650 | 2013-01-03 | Lab none | Installed | Twitchell Rice Interface LI-7500 | Istalled on the Twitchell Rice tower to replace AIU-0674 for repair of the condensation problem. |
LI-7500A 0065 | 2013-01-03 | Lab none | Installed | West Pond Gas Analyzer CO2/H2O | Installed on the West Pond tower with AIU-0666 to swap out sn 2182 for calibration and AIU-0397 for repair of the condensation problem. |
LI-7500A 75H-2180 | 2013-01-03 | Lab none | Installed | Twitchell Rice Gas Analyzer CO2/H2O | Installed on the Twitchell Rice tower with AIU-0614 to swap out sn 2176 for calibration and AIU-0674 for repair of the condensation problem. |
LI-7500A 75H-2182 | 2013-01-03 | West Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the West Pond tower for calibration and to swap out the 7550 box (sn 0397) for repair of the condensation problem. Replaced by sn 0065 and AIU-0666. |
LI-7500A 75H-2176 | 2013-01-03 | Twitchell Rice Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Twitchell Rice tower for calibration and to swap out the 7550 box (sn 0674) for repair of the condensation problem. Replaced by sn 2180 and AIU-0614. |
LI-7550 AIU-0666 | 2012-12-21 | Factory | Returned | Lab none | Returned from Licor with 75H-0065 after repair of condensation problem. The external connector board has been completely sealed and new firmware v6.0.0 has been loaded to the box. Pressure Sensor sn PX-2763: A0=57.595, A1=15.231 |
LI-7500A 0065 | 2012-12-21 | Lab none | Returned | Lab none | Returned from Licor with AIU-0666 after repair of condesation problem. Condesation was removed from the source lens and the detector was noisy so was also replaced. The 7500 was recalibrated.
CO2: A=1.57076E2, B=5.35482E3, C=5.64100E7, D=-1.93457E10, E=2.76431E12, XS=0.0025, Z=-9.00000E-4, zero=0.8901, span=1.0010 H2O: A=4.99824E3, B=6.61401E6, C=-1.74913E9, XS=-0.0007, Z=1.06000E-2, zero=0.7016, span=0.9935, CX=32000 SS: B=2.691, C=4.178 |
LI-7550 AIU-0650 | 2012-12-21 | Factory | Returned | Lab none | Returned from Licor with 75H-2180 after repair for the condensation problem. The external connector circuit board at the bottom of the box has been completely sealed. New firmware v6.0.0 has been loaded on the box and requires new Windows interface software. Pressure Sensor sn PX-2820 A0=57.729, A1=15.323. |